OpenStudy (anonymous):

Solve by using the quadratic formula. 3x^2 - 4x = 2.

OpenStudy (nurali):

The quadratic formula is given by: x = [-b ±√(b)² - 4ac / 2a] In the given equation: 3x² - 4x - 2 = 0 a=3, b=-4 and c=-2 So we punch all this into the formula: x = -(-4) ±√(-4)²- 4(3)(-2) / 2(3) x = 4 ±√16 + 24 / 2(3) x = 4/6 ± 1/6√40 x = 2/3 ±1/3√10